# Compression, Impact and Hot Rebound Flows from Coronal Rain Downflows

### v1

@@ Abstract
Understanding the processes associated with coronal rain due to the thermal non-equilibrium (TNE) and thermal instability (TI) scenario can help us understand coronal heating. We aim to study the properties of a quiescent coronal rain event and its effect on the solar atmosphere. We utilise space-based data from the High-Resolution Imager in Extreme Ultraviolet of Solar Orbiter, the Atmospheric Imaging Assembly of the Solar Dynamics Observatory, and the Slit-Jaw Imager (SJI) from the Interface Region Imaging Spectrograph from November 1st, 2023. During the coronal rain shower, the coronal loop exhibits large EUV variability and drastic changes in sub-structure. Coronal rain clumps with total velocities between 72 km s-1 and 87 km s-1 and cool EUV absorbing core sizes of $\approx$600 km and densities of $\approx 5\times 10^{11}$ cm-3 are seen to fall with a strong compression ahead. During the compression we measure a low polytropic index with $\gamma=1.085$, suggesting the presence of molecules. The rain shower carries a total of $3.09\times 10^{26}$ erg, and the clumps produce impacts seen in all EUV channels and in SJI 1400 Å. The impacts generate hot rebound flows with temperatures of $10^{6.2}-10^{6.3}$~K and velocities of 858785-87 km s-1, which refill and reheat the loop but carry less than $20\%$ of the clumps’ kinetic energies. We find signatures of a steady footpoint heating, in agreement with the TNE-TI scenario, with an estimated amplitude of $2.56\times 10^{-2}$~erg cm-3 s-1 in agreement with active region estimates. Coronal rain may therefore be a good proxy for the total integrated heating that gives birth to TNE-TI.
@@
pubyear: 2025pagerange: Compression, Impact and Hot Rebound Flows from Coronal Rain DownflowsReferences
@@ 1 Introduction
Until the 1940s, it was widely accepted that temperatures within the sun decrease radially outwards. However, this viewpoint was called into question and a discovery (Grotrian, 1934; Edlen, 1945) was made indicating that the solar corona reheats to temperatures exceeding millions of degrees Kelvin. The heating mechanism behind coronal heating has eluded scientists for many years to what is now known as the ‘coronal heating problem’. Through the help of countless observations and numerical simulations (Van Doorsselaere et al., 2020; van der Holst et al., 2014; Dere, 1994) it is evident that the heating mechanism involves the release of magnetic energy through processes such as magnetic reconnection and magnetohydrodynamic (MHD) waves.
@@
In addition to the descending coronal rain, the recent study by Antolin et al. (2023) identified an atmospheric response to the rain impact. They found speeds of the upward propagating features following the rain impact in the range of 5013050-130  km  s-1, and interpreted the lower and upper values in this range as corresponding to upflows and rebound shocks, respectively. This atmospheric response associated with coronal rain has also been predicted in numerical studies, e.g. by Müller et al. (2003).
@@
Neighboring coronal rain clumps are commonly observed to fall in tandem with each other, spanning over a wide cross-section of coronal loops. This phenomenon is known as a ‘rain shower’ (Antolin & Rouppe van der Voort, 2012; Şahin & Antolin, 2022). Local thermal instability (TI) within a coronal loop in a state of thermal-non-equilibrium (TNE) is said to be the cause of coronal rain, and is also thought to be the synchronising mechanism leading to rain showers. This is supported by numerical simulations by Fang et al. (2013), where a first TI-driven condensation perturbs neighboring critically stable field lines leading to a fast growth of the condensation in the direction perpendicular to the field. TNE occurs when there is strongly stratified and relatively long lasting heating. Namely, the loop footpoints need to be heated roughly ten times more than the loop apex, without much heating asymmetry between the two footpoints and the heating needs to be sustained longer than the radiative cooling time (Klimchuk & Luna, 2019; Johnston et al., 2019). After a rapid heating phase, the coronal loop becomes increasingly dense due to sustained chromospheric evaporation, which in turn leads to an over-density compared to that needed to achieve thermal equilibrium. Radiative losses then become larger than the input from thermal conduction and enthalpy, leading to a runaway cooling and localised TI in the corona. Condensations occur, which fall as coronal rain largely evacuating the loops. This TNE-TI scenario repeats if the heating persists, leading to TNE-TI cycles of heating and cooling that manifest as long-period EUV intensity pulsations (Auchère et al., 2014; Froment et al., 2015) and periodic coronal rain (Auchère et al., 2018).
However, there is still an ongoing debate on the role of TI, and the exact route to catastrophic cooling (Klimchuk, 2019; Waters & Stricklan, 2024). Furthermore, the overall prevalence of TNE-TI in the solar atmosphere is unknown. Şahin et al. (2023), conducted a statistical analysis of the ‘rain showers’ for one active region (AR) and found the coronal volume in a state of TNE-TI to be at least 50% the AR volume. Determining the coronal volume subject to TNE-TI is important because of its strong link to coronal heating. Indeed, the specific spatiotemporal distribution of heating required to produce coronal rain constitutes a strong constraint for heating mechanisms.
@@
Antolin et al. (2022) conducted 2.5D radiative MHD simulations which depict TI-driven catastrophic cooling leads to the formation of cool and dense cores at the head of the rain emitting in chromospheric lines. These cores are surrounded by thin but strongly emitting shells in TR lines that elongate in the wake of the rain, in what is known as the condensation corona transition region (CCTR).
There are three different types of coronal rain (Antolin & Froment, 2022). The first is the quiescent type usually observed in non-flaring ARs and does not rely on a specific magnetic topology. This study focuses on this type. The second is flare-driven coronal rain which is associated with the cool material observed in the gradual phase of flare loops. Lastly, there is a prominence-coronal rain hybrid, also known as coronal spider or cloud prominence, which involves a complex magnetic field with a fan-spine (null-point) topology at the top of loop arcades.
@@
With the help of the Atmospheric Imaging Assembly (AIA, Lemen et al., 2012) onboard the Solar Dynamics Observatory (SDO, Pesnell et al., 2012), alongside the Slit Jaw Imager (SJI) onboard the Interface Region Imaging Spectrograph (IRIS, De Pontieu et al., 2014), and the High-Resolution Imager of the Extreme Ultra Violet Instrument (HRIEUV) aboard the Solar Orbiter mission (Rochus et al., 2020; Müller et al., 2020), this study provides a detailed analysis of quiescent coronal rain observed during the SolO perihelion of November 1st 2023. Through the variety of wavelengths and viewpoints, we examined various properties of coronal rain. In particular, we focus on the compression produced by the rain and its effect on the coronal loop upon impact on the lower atmosphere. In section 2; we discuss our observation; 3; discuss our methodology; 4; discuss our findings; and a detailed analysis is discussed in 5.
@@ 2 Observations
HRIEUV was positioned at 0.56 AU from the Sun and captured images in a format of $2048\times 2048$ pixels. This provided a full field-of-view (FOV) of $1007.6"\times 1007.6$", with a pixel scale of 0.492 arcseconds per pixel and a pixel sampling of 199.83 km  per pixel. The observation began on the 1st of November 2023 at 00:43:41 UT, with the last image recorded at 01:43:31 UT. An extra 3.65 minutes was added to the start and end times to account for the time it takes light to reach the AIA compared to HRIEUV. This observation was part of the ‘R SMALL HRES MCAD AR-Heating’ Solar Orbiter Planning Sheet (SOOP) which provided a cadence of 10 s, an exposure time of 3 s and a total duration of approximately 60 min. In the 174 Å  passband, Fe ix (171 Å) and Fe x (174 Å) are the ions that dominate with a peak formation temperature of $10^{6}$~K. The images have been prepped with the euiprep routine to level 2, which reduces jitter and pointing errors. Significant jitter remained and a cross-correlation technique was applied to remove this, following Chitta et al. (2022).
@@
While most AIA passbands operated at a cadence of 12 s, AIA 1600 Å   and AIA 1700 Å  operated at a cadence of 24 seconds. AIA has a pixel scale of 0.6 arcseconds per pixel. The AIA dataset obtained for this study starts on the 1st of November at 00:29:26 UT and ends at 01:59:40 UT allowing for a total duration of roughly 90 min. The pixel sampling obtained from AIA is 431 km  per pixel.
@@
We use SJI 1400 Å  data from IRIS, which predominantly captures the Si iv resonance lines at 1394 Å   and 1402 Å, which originate from plasma characteristic of the transition region and have a mean formation temperature of 0.08 MK. The SJI 1400  Å  operated with a cadence of 10.5 s and an exposure time of 4 s. Each image obtained by SJI 1400 Å  has dimensions of $1001\times 1093$ pixels with a pixel scale of 0.16635 arcseconds per pixel. This provided a full field of view of $166.52"\times 181.82$" and a pixel sampling size of 120  km  per pixel. The FOV centre is -131.5 WEST and 199.24 North. The observation began on the 1st of November at 00:38:17 UTC and concluded at 01:49:49 UTC allowing for a total duration of approximately 70 min. IRIS conducted a very large sparse 2-step raster observation for this event. The slit did not cross the event and IRIS only party captured the loop footpoint as shown in Figure 2.
@@ 3 Methodology
$\text{DEM}(T)dT=\int_{0}^{\infty}n_{e}n_{p}(T)dl,$ (1)
where $n_{e}$ and $n_{p}$ are the electron and proton number densities, respectively, and $l$ denotes the distance along the LOS. The total Emission Measure (EM) is the integration of the DEM over the temperature range:
@@
$\langle n_{e}\rangle_{\text{DEM}}=\sqrt{\frac{1.2\int_{T}\text{DEM}(T)dT}{l}}.$ (4)
@@
The DEM is calculated in IDL using the CHIANTI (version 10.1 developed by Dere et al. (2023)) atomic database, first introduced by Dere et al. (1997). The method used is the regularization method discussed in Hannah & Kontar (2012). We selected 16 temperature bins, ranging from $\log_{10}T=5.5$ to $\log_{10}T=7.05$ with a bin size of 0.1 in logarithmic space.
@@ 4.1.2 Loop Structure
Figure 9: Time-distance diagrams of the two paths shown in Figure 4. The cyan arrow in ‘Path 1’ points to a coronal strand that brightens prior to the rain falling. The black arrows highlight the EUV variability as bright strands appear to fade away and new strands emerge. The green arrow in ‘Path 2’ points to an EUV absorption feature caused by the rain clump.
@@ 4.2.1 Morphology and EUV variability
By analysing the absorption created by the rain in Figure 10 we estimated the length of downflow 1 to be $1.1\pm 0.3$ Mm. This length could be much longer Antolin & Rouppe van der Voort (as seen in previous rain observations, e.g. 2012), but not producing enough absorption to be observed. However, we believe that downflow 1 is a collection of smaller, unresolved clumps (as seen in AIA 304 Å) and to obtain a more accurate length, we multiplied the total velocity over the time between downflow 1 and the beginning of the upward flow (equal to 120 s, which matches the time interval where most of the enhanced brightening is observed in the impact region). This leads to a length of $8.7\pm 1.2$ Mm.
@@ 4.3 Compression by the rain
We now analyse the compression produced by the rain during the fall and show that the ratio of specific heats (and polytropic index) can be estimated from it. For this, we first calculate the electron number densities and DEM-weighted temperatures in the compressed plasma. Using the methods described in Section 3 we calculate these quantities along the blue and red lines in Figure 14, which follow, respectively, the compression produced by the rain and the parallel to this at a time 10 min prior, when the rain compression is not observed (but the conditions in the loop are otherwise assumed to be very similar). We denote the latter conditions (from the red line) as the ‘loop’ conditions (without rain compression). The result is shown in blue in Figure 15.
Because of the relatively low efficiency of thermal conduction in the compressed region (the energy cannot be conducted upwards due to the presence of the rain acting like a cork), and a low loss through radiation due to the relatively small time interval, we can assume the compression to be adiabatic, which implies:
@@
where $T_{c}$ and $T_{l}$ denote, respectively, the compression and loop temperatures obtained from the DEM-weighted temperature. Similarly, $n_{e,c}$ and $n_{e,l}$ denote the electron number densities of the compression and of the loop, respectively, obtained from equation 4. $\gamma$ denotes the ratio of specific heats. We then have:
@@
To calculate the electron number density we took $l$ in Equation 4, as the width of the observed compression. We obtained a compression width of $0.557\pm 0.01$ Mm observed in HRIEUV and $0.690\pm 0.01$ Mm observed in AIA 171 Å. The larger compression width in AIA is likely due to the instrument’s lower spatial resolution. Therefore, we used for $l$ the width obtained in HRIEUV.
For the compression values of density and temperature, we choose the point of maximum compression seen in Figure 15. Using this location in time and space, we take the number density of the loop at the same height but 10 minutes before (along the red line in Figure 14). This leads to a value of $T_{c}=(1.04\pm 0.05)\times 10^{6}~\text{K}$, $n_{e,c}=(5.59\pm 0.19)\times 10^{9}~\text{cm}^{-3}$, $T_{l}=(1.00\pm 0.09)\times 10^{6}~\text{K}$, $n_{e,l}=(3.43\pm 0.16)\times 10^{9}~\text{cm}^{-3}$. We then obtain $\gamma=(1.085\pm 0.218)$, reflecting the isothermal behaviour during the compression seen in Figure 15 and the reionisation of the neutral plasma in the rain.

### v2

@@ Abstract
Studying coronal rain formation through thermal non-equilibrium (TNE) and thermal instability (TI) provides insights into coronal heating mechanisms. We analysed a quiescent coronal rain event using space-based observations from the High-Resolution Imager in Extreme Ultraviolet (HRIEUV) of Solar Orbiter (SolO), the Atmospheric Imaging Assembly (AIA) of the Solar Dynamics Observatory (SDO), and the Slit-Jaw Imager (SJI) from the Interface Region Imaging Spectrograph (IRIS) from November 1st, 2023. During the coronal rain shower, the coronal loop exhibits substantial EUV variability and structural changes. Rain clumps fell at 728772-87 km s-1 with cool EUV absorbing core sizes of $\approx$600 km and densities of $\approx 6\times 10^{11}$ cm-3 preceded by strong compressions. These mostly isothermal compressions suggest energy transfer into the rain, decelerating it and possibly reducing cooling rates – consistent with accretion braking timescales. The shower carried microflare-level energy ($4.64\times 10^{26}$ erg), with clumps producing impacts that reach the lower transition region and are visible across all EUV channels and in SJI 1400 Å. The impacts generated hot rebound flows ($10^{6.2}-10^{6.3}$~K, 858785-87 km s-1) that refilled and reheated the loop but carried less than $15\%$ of the clumps’ kinetic energy. We detected steady footpoint heating signatures consistent with the TNE-TI scenario, with an estimated amplitude of $10^{-2\pm 0.3}$~ erg cm-3  s-1 and heating scale heights of 2102-10 Mm, matching active region values. Coronal rain may thus serve as both a template for accretion braking and a proxy for integrated heating driving TNE-TI cycles.
@@
pubyear: 2026pagerange: Compression, Impact and Hot Rebound Flows from Coronal Rain DownflowsReferences
@@ 1 Introduction
The recent study by Antolin et al. (2023) identified an atmospheric response to the rain impact. They found speeds of the upward propagating features following the rain impact in the range of 5013050-130  km  s-1, and interpreted the lower and upper values in this range as corresponding to upflows and rebound shocks, respectively. This atmospheric response associated with coronal rain has also been predicted in numerical studies, e.g. by Müller et al. (2003).
@@
Neighboring coronal rain clumps are commonly observed to fall in tandem with each other, spanning over a wide cross-section of coronal loops. This phenomenon is known as a ‘rain shower’ (Antolin and Rouppe van der Voort, 2012; Şahin and Antolin, 2022). Local thermal instability (TI) within a coronal loop in a state of thermal-non-equilibrium (TNE) is said to be the cause of coronal rain, and is also thought to be the synchronising mechanism leading to rain showers. This is supported by numerical simulations by Fang et al. (2013). TNE occurs when there is strongly stratified and relatively long-lasting heating. Namely, the loop footpoints need to be heated roughly ten times more than the loop apex, without much heating asymmetry between the two footpoints and the heating needs to be sustained longer than the radiative cooling time (Klimchuk and Luna, 2019; Johnston et al., 2019). After a rapid heating phase, the coronal loop becomes increasingly dense due to sustained chromospheric evaporation, which in turn leads to an over-density compared to that needed to achieve thermal equilibrium. Radiative losses then become larger than the input from thermal conduction and enthalpy, leading to a runaway cooling and localised TI in the corona (Antolin, 2020; Keppens et al., 2025). Condensations occur, which fall as coronal rain largely evacuating the loops. This TNE-TI scenario repeats if the heating persists, leading to TNE-TI cycles of heating and cooling that manifest as long-period EUV intensity pulsations (Auchère et al., 2014; Froment et al., 2015) and periodic coronal rain (Auchère et al., 2018).
@@
Antolin et al. (2022) conducted 2.5D radiative MHD simulations which depict TI-driven catastrophic cooling that leads to the formation of cool and dense cores at the head of the rain emitting in chromospheric lines. These cores are surrounded by thin but strongly emitting shells in TR lines (the CCTR) that elongate in the wake of the rain.
@@
This work focuses on the coronal rain observed in non-flaring active regions (ARs), known as quiescent coronal rain. With the help of the Atmospheric Imaging Assembly (AIA, Lemen et al., 2012) onboard the Solar Dynamics Observatory (SDO, Pesnell et al., 2012), alongside the Slit-Jaw Imager (SJI) onboard the Interface Region Imaging Spectrograph (IRIS, De Pontieu et al., 2014), and the High-Resolution Imager of the Extreme Ultra Violet Instrument (HRIEUV) aboard the Solar Orbiter (SolO) mission (Rochus et al., 2020; Müller et al., 2020), this study provides a detailed analysis of quiescent coronal rain observed during the SolO perihelion of November 1st 2023. Through the variety of wavelengths and viewpoints, we examined various properties of coronal rain. In particular, we focus on the compression produced by the rain and its effect on the coronal loop upon impact on the lower atmosphere. In section 2 we discuss our observation, in section 3 we discuss our methodology, in section 4 we discuss our findings, and a detailed analysis is discussed in section 5.
@@ 2 Observations
HRIEUV was positioned at 0.56 AU from the Sun and captured images in a format of $2048\times 2048$ pixels. This provided a full field-of-view (FOV) of $1007\arcsec.6\times 1007\arcsec.6$, with a pixel scale of 0″.492 per pixel and a pixel sampling of 199.83 km  per pixel. The observation began on the 1st of November 2023 at 00:43:41 UT, with the last image recorded at 01:43:31 UT. An extra 3.65 minutes was added to the start and end times to account for the time it takes light to reach the AIA compared to HRIEUV. This observation was part of the ‘R SMALL HRES MCAD AR-Heating’ Solar Orbiter Planning Sheet (SOOP) which provided a cadence of 10 s, an exposure time of 3 s and a total duration of approximately 60 min. In the 174 Å  passband, Fe ix (171 Å) and Fe x (174 Å) are the ions that dominate with a peak formation temperature of $10^{6}$~K. The images have been prepped with the euiprep routine to level 2, which reduces jitter and pointing errors. Significant jitter remained and a cross-correlation technique was applied to remove this, following Chitta et al. (2022). We have estimated the accuracy of this algorithm by randomly adding or subtracting a 1-$\sigma$ errors (photon shot noise) to each pixel and running the cross-correlation algorithm again. Repeating this process 100 times has yielded a standard deviation on the displacement shifts on the order of 10 km.
@@
While most AIA passbands operated at a cadence of 12 s, AIA 1600 Å   and AIA 1700 Å  operated at a cadence of 24 seconds. AIA has a pixel scale of 0″.6 per pixel. The AIA dataset obtained for this study starts on the 1st of November at 00:29:26 UT and ends at 01:59:40 UT allowing for a total duration of roughly 90 min. The pixel sampling obtained from AIA is 431 km  per pixel.
@@
We use SJI 1400 Å (OBS-ID 3440007611), data from IRIS, which predominantly captures the Si iv resonance lines at 1394 Å   and 1402 Å, which originate from plasma characteristic of the transition region and have a mean formation temperature of 0.08 MK. The SJI 1400  Å  operated with a cadence of 10.5 s and an exposure time of 4 s. Each image obtained by SJI 1400 Å  has dimensions of $1001\times 1093$ pixels with a pixel scale of 0″.166 per pixel. This provided a full field of view of $166\arcsec.52\times 181\arcsec.82$ and a pixel sampling size of 120  km  per pixel. The FOV centre is -131.5 West and 199.24 North. The observation began on the 1st of November at 00:38:17 UTC and concluded at 01:49:49 UTC allowing for a total duration of approximately 70 min. IRIS conducted a very large sparse 2-step raster observation for this event. The slit did not cross the event and IRIS only party captured the loop footpoint as shown in Figure 2.
@@ 3 Methodology
$\text{DEM}(T)dT=\int_{0}^{\infty}n_{e}n_{p}(T)d\ell,$ (1)
where $n_{e}$ and $n_{p}$ are the electron and proton number densities, respectively, and $\ell$ denotes the distance along the LOS. The total Emission Measure (EM) is the integration of the DEM over the temperature range:
@@
$\langle n_{e}\rangle_{\text{DEM}}=\sqrt{\frac{1.2\int_{T}\text{DEM}(T)dT}{\ell}}.$ (4)
@@
The DEM is calculated in IDL using the CHIANTI (version 10.1 developed by Dere et al. (2023)) atomic database, first introduced by Dere et al. (1997). The method used is the regularization method discussed in Hannah and Kontar (2012). We selected 16 temperature bins, ranging from $\log_{10}T=5.5$ to $\log_{10}T=7.05$ with a bin size of 0.1 in logarithmic space.
@@ 4.1.2 Loop Structure
Figure 9: Time-distance diagrams of the two paths shown in Figure 4. Where 0 Mm corresponds to the right side of Path 1 and 2 in Fig. 4. The cyan arrow in ‘Path 1’ points to a coronal strand that brightens prior to the rain falling. The black arrows highlight the EUV variability as bright strands appear to fade away and new strands emerge. The green arrow in ‘Path 2’ points to an EUV absorption feature caused by the rain clump. Note that it is preceded by a brightening feature, which corresponds to the compression analysed in section 4.3.
@@ 4.2.1 Morphology and EUV variability
By analysing the absorption created by the rain in Figure 10 we estimated the length of downflow 1 to be $1.1\pm 0.3$ Mm. This length could be much longer (as seen in previous rain observations, e.g. Antolin and Rouppe van der Voort, 2012), but not producing enough absorption to be observed. However, we believe that downflow 1 is a collection of smaller, unresolved clumps (as seen in AIA 304 Å) and to obtain a more accurate length, we infer it by multiplying the total velocity over the time between downflow 1 and the beginning of the upward flow (equal to $120\pm 30$ s, which matches the time interval where most of the enhanced brightening is observed in the impact region). This leads to a length of $8.7\pm 0.3$ Mm.
We note that along all of the splines plotted we can detect traces of the rain falling in the corresponding time-distance diagrams, indicating that the rain shower has a width at least equal to the transverse extent of the splines, which is 1 Mm.
@@ 4.3 Compression by the rain
We now analyse the compression produced by the rain during the fall and show that the ratio of specific heats (and polytropic index) can be estimated from it. For this, we first calculate the electron number densities and DEM-weighted temperatures in the compressed plasma. Using the methods described in Section 3 we calculate these quantities along the blue and red lines in Figure 14, which follow, respectively, the compression produced by the rain (subindex $c$) and the parallel to this at a time 10 min prior, when the rain compression is not observed (but the conditions in the loop are otherwise assumed to be very similar). We denote the latter conditions (from the red line) as the ‘loop’ conditions (without rain compression, with subindex $l$). The result is shown in blue in Figure 15.
To calculate the electron number density we took $\ell$ in Equation 4, as the width of the observed compression. We obtained a compression width of $0.557\pm 0.01$ Mm observed in HRIEUV and $0.690\pm 0.01$ Mm observed in AIA 171 Å. The larger compression width in AIA is likely due to the instrument’s lower spatial resolution. Therefore, we used for $\ell$ the width obtained in HRIEUV. From Figure 15 we see that the compression is to a large extent isothermal. At maximum compression we have $n_{e,c}=(5.59\pm 0.19)\times 10^{9}~\text{cm}^{-3}$, $n_{e,l}=(3.43\pm 0.16)\times 10^{9}~\text{cm}^{-3}$, $T_{c}=(1.04\pm 0.05)\times 10^{6}~\text{K}$, and $T_{l}=(1.00\pm 0.09)\times 10^{6}~\text{K}$. That is, the density increases by a factor of 1.6, while the temperature only changes by a factor of 1.04 (i.e. 40,000 K).
To understand this it is important to note that the rain acts like a cork for the transfer of energy upwards via thermal conduction. However, it can still be radiated in-situ in the compressed region or transferred downwards to the transition region. We estimate the conduction timescale $\tau_{\text{cond}}$, the radiative timescale $\tau_{\text{rad}}$ and the advective timescale $\tau_{\text{adv}}$, which is the time to impact of the rain. We have:
$\tau_{\text{cond}}$ $=$ $\frac{3}{2}\frac{n_{e,c}L^{2}k_{B}}{10^{-6}T_{c}^{5/2}}\approx 300~\text{s}$ (9)
$\tau_{\text{rad}}$ $=$ $\frac{3}{2}\frac{k_{B}T}{n_{e,c}\Lambda}\approx 120-170\text{~s}$ (10)
$\tau_{\text{adv}}$ $=$ $\frac{L}{v}\approx 70~\text{s},$ (11)
where we have taken L=5000L=5000~km as the length over which the compression is observed.
We can see that the condition of adiabaticity $\tau_{\text{adv}}<\tau_{\text{cond}},\tau_{\text{rad}}$ holds to some extent. If so, we have:
@@
where $\gamma$ denotes the ratio of specific heats. We then have:
@@
Using the values at maximum compression stated previously, we obtain $\gamma=(1.085\pm 0.218)$, reflecting the isothermal behaviour during the compression seen in Figure 15. However, it is possible that the density of the compressed region just underneath the rain is actually not properly resolved, and much denser, in which case the radiative losses could be the main sink of energy.